# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
Figure 4: Long coherence times and high-fidelity single-qubit gates in a large atom array. a, Array-averaged Rabi oscillations between the hyperfine clock states $|0\rangle$ and $|1\rangle$. The fitted Rabi frequency is 24.611(1) kHz. The observed decay after several hundred microseconds stems from the spatially-varying Rabi frequency (Ext. Data Fig. 5c). b, Array-averaged Ramsey oscillations. During free evolution, the microwave drive field is detuned by 1 kHz, resulting in Ramsey oscillations. The characteristic decay time of these oscillations is $T_{2}^{*}=14.0(1)\,\mathrm{ms}$ fitting the average signal of all atoms. The light blue dashed line shows the decay time $T_{2}^{*(\mathrm{site})}=25.5\ \mathrm{ms}$ from fitting individual sites first and averaging the decay time afterwards. c, Measurement of the dephasing time $T_{2}$ after dynamical decoupling. After an initial $\pi/2$ pulse, a variable number of XY16 dynamical decoupling cycles with a fixed time $\tau=12.5\ \mathrm{ms}$ between $\pi$ pulses are used to offset the reversible dephasing. The phase of the final $\pi/2$ pulse is chosen to be either 0 or $\pi$, and subtracting the population difference in these two cases provides the coherence contrast. The contrast decay is fitted to obtain $T_{2}=12.6(1)\,\mathrm{s}$. d, Randomized benchmarking of the global single-qubit gate fidelity. For each number of Clifford gates, 60 different random gate strings of this length are applied, after which the overall inverse of the string is applied. The translucent markers represent the return probability for each string of gates, while the solid green markers indicate the averaged return probability over the 60 different strings. The decay of the final population to $1/2$ is fitted to $(1-d)^{N}$ and $F_{c}=1-d/2$ represents the average Clifford gate fidelity.
@@
Moreover, we accurately characterize the imaging survival probability, without assuming any parameters, by performing 80 ms repeated imaging up to 1000 times, after which ${\sim}90$% of initially loaded atoms still survive (Fig. 3b). This corresponds to a steady-state imaging survival probability of $99.98952(1)\%$, mostly limited by vacuum lifetime. This, to the best of our knowledge, surpasses prior studies reporting record steady-state imaging survival using single alkaline-earth metal covey_2000-times_2019 and alkali-metal blodgett_imaging_parsed_2023 atoms in optical tweezers. These results, and the uniformity of imaging survival across the array (Ext. Data Fig. 4a), enable low-loss, high-fidelity detection of single atoms in large-scale arrays, crucial components for the practical use of the system (see Discussion and Outlook).
@@
Key to recent progress in quantum computing and metrology with neutral atoms is the ability to encode a qubit in long-lived states of an atom, such as hyperfine states bluvstein_quantum_2022, graham_multi-qubit_2022, nuclear spin states barnes_assembly_2022, huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, or optical clock states norcia_seconds-scale_2019, madjarov_atomic-array_2019. In cesium atoms, the hyperfine ground states ($|F=3,\ m_{F}=0\rangle\equiv|0\rangle$ and $|F=4,\ m_{F}=0\rangle\equiv|1\rangle$) provide such a subspace for storing quantum information. Furthermore, entanglement via Rydberg interactions can be readily transferred to this qubit to realize high-fidelity two-qubit gates evered_high-fidelity_2023, finkelstein_universal_2024. We now demonstrate the storage and manipulation of quantum information in a large-scale atom array by measuring the coherence time and global single-qubit gate fidelity. To this end we use a microwave horn to drive the hyperfine transition (Fig. 4a).
@@
Preserving the coherence of a quantum system as it is scaled up is a known challenge across platforms for quantum computing and simulation kjaergaard_superconducting_2020. This difficulty persists even for neutral atoms, albeit at a lower level, due to residual interactions with a noisy and inhomogeneous electromagnetic environment, particularly with tweezers themselves (due to induced scattering and dephasing). Using far-off-resonant optical tweezers helps to preserve coherence, since the differential light shift of the hyperfine qubit decreases as $1/\Delta_{\mathrm{tweezer}}$ at constant trap depth, and the scattering rate as $1/\Delta_{\mathrm{tweezer}}^{3}$, where $\Delta_{\mathrm{tweezer}}$ is the tweezer laser detuning relative to the dominant electronic transition kuhr_analysis_2005, ozeri_hyperfine_2005, ozeri_errors_2007. We verify this by measuring a depolarization time of $T_{1}=119(1)$ s (Ext. Data Fig. 5e), and an ensemble dephasing time in the atom array of $T_{2}^{*}=14.0(1)$ ms (Fig. 4b), limited by trap depth inhomogeneity. Measured site-by-site, the dephasing time is $T_{2}^{*(\mathrm{site})}=25.5$ ms, consistent with being limited by an atomic temperature of ${\sim}4.8\ \mathrm{\mu K}$ during microwave operation kuhr_analysis_2005. We also measure the standard deviation of the qubit frequency in the array to be $1.5\times 10^{-9}$, or 14 Hz in absolute value (Ext. Data Fig. 5d).
@@
The remaining reversible dephasing can be further mitigated by dynamical decoupling. By applying cycles of XY16 sequences gullion_new_1990, souza_robust_2012 with a dwell time of 12.5 ms between $\pi$ pulses, the measured dephasing time is $T_{2}=12.6(1)$ s, thereby setting a new benchmark for the coherence time of an array of hyperfine qubits in optical tweezers bluvstein_quantum_2022, graham_multi-qubit_2022 (Fig. 4c). Combined with recent advances in neutral atom quantum processors based on qubit transport beugnon_two-dimensional_2007, bluvstein_quantum_2022, bluvstein_logical_2024, which require dozens of microseconds for a set of parallel gates (and thus long coherence times), such long coherence times could enable the realization of complex quantum algorithms with thousands of qubits in the near-term.
@@
Discussion and outlook
We have demonstrated scaling of the optical tweezer array platform to over 6,100 individually-trapped atoms. While some crucial ingredients have yet to be demonstrated, especially the ability to rearrange atoms and entangle atomic qubits, we show equal or superior results to state-of-the-art experiments on several key metrics that are known bottlenecks for quantum computing, quantum simulation and quantum metrology. Notably, we achieve record-high imaging survival covey_2000-times_2019 alongside competitive imaging fidelity. We measure record room-temperature vacuum-limited lifetime covey_2000-times_2019 and coherence time in alkali metal atom tweezer arrays bluvstein_quantum_2022, graham_multi-qubit_2022, with a high global single-qubit gate fidelity, only limited by technical noise. We thereby largely circumvent aspects of degradation in classical and quantum control commonly faced by platforms as they are scaled up.
@@
Our results usher in a new generation of neutral atom quantum processors based on several thousand of qubits, as well as large-scale programmable devices enabling advances in quantum metrology rosenband_exponential_2013, ludlow_optical_2015, norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024 and simulation haferkamp_closing_2020, orourke_entanglement_2023, julia-farre_amorphous_2024. For example, our platform — with the demonstrated qubit numbers — could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2023. Tweezer clocks could be scaled using near-infrared, high-power tweezers for loading and imaging tao_high-fidelity_2023 before transferring atoms to magic-wavelength traps for clock operation norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024. We also foresee applications in quantum simulation for problems where boundary effects play an important role bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, orourke_entanglement_2023, which can be minimized with the large system sizes demonstrated here. In order to realize directions requiring interactions between atoms, we plan on exciting to Rydberg states saffman_quantum_2010. Promisingly, cesium provides a two-photon excitation pathway through the $7\textrm{P}_{3/2}$ intermediate state with a 455 nm lower photon transition and 1060 nm upper photon transition. With a large detuning from the intermediate state to minimize scattering, and with widely available laser powers at these wavelengths, it is possible to achieve Rabi frequencies as high as 5 MHz with minimal inhomogeneity across the full field of view.
Concerning universal quantum computation, the unusually large 1.5 mm field of view (FOV) of our objective provides enough space to engineer zones around our current array (900 $\mathrm{\mu m}$ in diameter) for interaction, storage, qubit rotation, and mid-circuit readout, enabling operations on qubits without disturbing adjacent atoms beugnon_two-dimensional_2007, bluvstein_quantum_2022. In order to utilize such a zone architecture for quantum computation, it is imperative to dynamically rearrange atoms across the optical tweezer array barredo_atom-by-atom_2016, endres_atom-by-atom_2016. Using commercially available acousto-optic deflectors with large field-of-view (${>}500\,\mathrm{\mu m}$ with our optical parameters) and fast steering mirrors, we foresee a straightforward pathway to low-loss transport of atoms across the complete objective field-of-view of 1.5 mm. This could be combined with recent technical developments that allow for high-speed parallel addressing of 10,000 sites zhang_scaled_2024, and demonstrations of ancilla-based mid-circuit readout techniques singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024, to ultimately pave the way for repeated quantum error correction in neutral atom arrays with thousands of physical qubits.
For most quantum simulation and computation tasks, rearrangement of the stochastically loaded atom array into a defect-free array is required barredo_atom-by-atom_2016, endres_atom-by-atom_2016. The probability of obtaining a defect-free array can be limited by several factors, including finite imaging fidelity, imaging survival, imperfections in the rearrangement sequence, or finite vacuum-limited lifetime. We can provide an early estimate assuming a rearrangement time much less than our vacuum lifetime, that based on imaging survival alone, the probability to obtain a defect-free array of  6,040 rearranged atoms can be up to 39.5%, with optimal choices of imaging thresholding for this purpose (Ext. Data Fig. 4b). The average total number of defects corresponding to this estimation is 0.93.
@@
Finally, our work indicates that further scaling of the optical tweezer array platform to tens of thousands of trapped atoms should be achievable with current technology, while essentially preserving high-fidelity control. In our present apparatus, several factors limit the number of sites. One limitation is the finite number of pixels of each SLM (which reduces the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers. By using available higher-resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently. Furthermore, we observe worsening optical aberrations at tweezer powers greater than that in the present study due to thermal heating of the objective. This is the main limitation on atom number for the results in this work, even after aberrations were mitigated using the SLM (Methods). This constraint could be circumvented in the short term by dissipating heat from the objective more effectively to suppress detrimental aberrations and in the long term by utilizing an objective with a housing material that retains less heat; a combination of these upgrades should allow us to almost double the number of tweezers that we create using two fiber amplifiers. We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques. Atom numbers may further be increased in our array with the same number of tweezers by utilizing enhanced loading brown_gray-molasses_2019 or re-loading techniques shaw_dark-state_2023. Already in the near-term, we expect to increase the number of atomic qubits to over ten thousand with the current system using a subset of these techniques.

### v3

@@ Abstract
Figure 4: Long coherence times and high-fidelity single-qubit gates in a large atom array. a, Array-averaged Rabi oscillations between the hyperfine clock states $|0\rangle$ and $|1\rangle$. The fitted Rabi frequency is 24.611(1) kHz. The observed decay after several hundred microseconds stems from the spatially-varying Rabi frequency (Ext. Data Fig. 7b). b, Array-averaged Ramsey oscillations. During free evolution, the microwave drive field is detuned by 1 kHz, resulting in Ramsey oscillations. The characteristic decay time of these oscillations is $T_{2}^{*}=14.0(1)\,\mathrm{ms}$ fitting the average signal of all atoms. The light blue dashed line shows the decay time $T_{2}^{*(\mathrm{site})}=25.5\ \mathrm{ms}$ from fitting individual sites first and averaging the decay time afterwards. c, Measurement of the dephasing time $T_{2}$ after dynamical decoupling. After an initial $\pi/2$ pulse, a variable number of XY16 dynamical decoupling cycles with a fixed time $\tau=12.5\ \mathrm{ms}$ between $\pi$ pulses are used to offset the reversible dephasing. The phase of the final $\pi/2$ pulse is chosen to be either 0 or $\pi$, and subtracting the population difference in these two cases provides the coherence contrast. The contrast decay is fitted to obtain $T_{2}=12.6(1)\,\mathrm{s}$. d, Randomized benchmarking of the global single-qubit gate fidelity. For each number of Clifford gates, 60 different random gate strings of this length are applied, after which the overall inverse of the string is applied. For a given gate string length, each translucent marker of a given color represents the return probability for a string of gates while the solid green markers indicate the averaged return probability over the 60 different strings. The inset lists all the colors used to indicate the 60 random gate strings for a given length. The decay of the final population to $1/2$ is fitted to $(1-d)^{N}$ and $F_{c}=1-d/2$ represents the average Clifford gate fidelity.
@@
Moreover, we accurately characterize the imaging survival probability, without assuming any parameters, by performing 80 ms repeated imaging up to 1,000 times, after which ${\sim}90$% of initially loaded atoms still survive (Fig. 3b). This corresponds to a steady-state imaging survival probability of $99.98952(1)\%$, mostly limited by vacuum lifetime. This, to the best of our knowledge, surpasses prior studies reporting record steady-state imaging survival using single alkaline-earth metal covey_2000-times_2019 and alkali-metal blodgett_imaging_parsed_2023 atoms in optical tweezers. These results, and the uniformity of imaging survival across the array (Ext. Data Fig. 5a), enable low-loss, high-fidelity detection of single atoms in large-scale arrays, crucial components for the practical use of the system.
@@
Key to recent progress in quantum computing and metrology with neutral atoms is the ability to encode a qubit in long-lived states of an atom, such as hyperfine states bluvstein_quantum_2022, graham_multi-qubit_2022, nuclear spin states barnes_assembly_2022, huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, or optical clock states norcia_seconds-scale_2019, madjarov_atomic-array_2019. In cesium atoms, the hyperfine ground states ($|F=3,\ m_{F}=0\rangle\equiv|0\rangle$ and $|F=4,\ m_{F}=0\rangle\equiv|1\rangle$) provide such a subspace for storing quantum information (see Methods for state preparation and readout procedures). Furthermore, entanglement via Rydberg interactions can be readily transferred to this qubit to realize high-fidelity two-qubit gates evered_high-fidelity_2023, finkelstein_universal_2024. We now demonstrate the storage and manipulation of quantum information in a large-scale atom array by measuring the coherence time and global single-qubit gate fidelity. To this end we use a microwave horn to drive the hyperfine transition (Fig. 4a). For microwave operation we adiabatically ramp down tweezers to a depth of $k_{B}\times 55\ \mathrm{\upmu K}$.
@@
Preserving the coherence of a quantum system as it is scaled up is a known challenge across platforms for quantum computing and simulation kjaergaard_superconducting_2020. This difficulty persists even for neutral atoms, albeit at a lower level, due to residual interactions with a noisy and inhomogeneous electromagnetic environment, particularly with tweezers themselves (due to induced scattering and dephasing). Using far-off-resonant optical tweezers helps to preserve coherence, since the differential light shift of the hyperfine qubit decreases as $1/\Delta_{\mathrm{tweezer}}$ at constant trap depth, and the scattering rate as $1/\Delta_{\mathrm{tweezer}}^{3}$, where $\Delta_{\mathrm{tweezer}}$ is the tweezer laser detuning relative to the dominant electronic transition kuhr_analysis_2005, ozeri_hyperfine_2005, ozeri_errors_2007. We verify this by measuring a depolarization time of $T_{1}=119(1)$ s (Ext. Data Fig. 7d), and an ensemble dephasing time in the atom array of $T_{2}^{*}=14.0(1)$ ms (Fig. 4b), limited by trap depth inhomogeneity. Measured site-by-site, the dephasing time is $T_{2}^{*(\mathrm{site})}=25.5$ ms, consistent with being limited by an atomic temperature of ${\sim}4.3\ \mathrm{\upmu K}$ during microwave operation kuhr_analysis_2005. In Ext. Data Fig. 8a and in the Methods we present and discuss site-resolved qubit coherence data.
@@
The remaining reversible dephasing can be further mitigated by dynamical decoupling. By applying cycles of XY16 sequences gullion_new_1990, souza_robust_2012 with a period of 12.5 ms between $\pi$ pulses, the measured dephasing time is $T_{2}=12.6(1)$ s, thereby setting a new benchmark for the coherence time of an array of hyperfine qubits in optical tweezers bluvstein_quantum_2022, graham_multi-qubit_2022 (Fig. 4c). In addition, we investigate in Ext. Data Fig. 7g the coherence time at different trap depths, yielding notably $T_{2}=3.19(5)$ s at the full trap depth of $k_{B}\times 0.18$ mK. Although significantly lower, this result also surpasses previous known results with hyperfine qubits in a tweezer array. In the next section, we leverage the long coherence times demonstrated here to realize high-fidelity qubit transport beugnon_two-dimensional_2007 — the linchpin of zone-based state-of-the-art neutral atom quantum computation bluvstein_quantum_2022, bluvstein_logical_2024, finkelstein_universal_2024 — across record distances.
@@
Towards quantum computing with a zone-based architecture — coherent transport
We now focus more specifically on the practical implementation of a quantum computer, as it is a flagship application of our work and because it demands the most sophisticated toolbox of aforementioned use cases. Universal quantum computing requires single-qubit and two-qubit gates which can be implemented either through single-site addressing graham_multi-qubit_2022 or a zone-based architecture bluvstein_quantum_2022, bluvstein_logical_2024.
Zone-based architectures leverage the ability to dynamically move atoms in a coherence-preserving manner beugnon_two-dimensional_2007, bluvstein_quantum_2022, enabling long-range, non-local connectivity, which allows for less stringent quantum error correction bounds than grid-like connectivity bravyi_tradeoffs_2010, presenting a promising pathway to large-scale fault-tolerant quantum computation. A zone-based architecture also provides a path for mid-circuit readout bluvstein_logical_2024.
We depict a possible zone layout in Fig. 5a, which includes a storage zone large enough for more than 6,100 atoms. We do not foresee challenges in creating the zones themselves, e.g., Rydberg-based two-qubit gates should be feasible in a large interaction zone for more than 500 gates in parallel with state-of-the-art fidelities (SI §IV).
Figure 5: Long-distance and high-fidelity coherent transport. a, Proposed layout of a zone-based universal quantum processor with 6,100 atoms. Atoms anywhere in the storage zone can be transported with AODs to the interaction or readout zones in under 500 $\upmu$m. b, Schematic and atom survival for a single diagonal (blue) or straight (pink) move. Despite being shorter, a high-survival straight move needs to be executed more slowly than a diagonal one due to cylindrical lensing. The transport tweezer depth is $k_{B}\times 0.28$ mK. The following subfigures utilize diagonal transport. c, Coherence of an atom after being transported diagonally 610 $\upmu$m (blue) or simply held (grey). A single XY4 dynamical decoupling sequence is inserted to suppress dephasing. The move duration is 1.6 ms. d, Diagram of the sequence used to benchmark the move fidelity. Random Clifford gates are interleaved between each of the $M(<N)$ moves, and $N-M$ gates are added at the end to keep the total number of gates $N$ constant. e, f, g, Results of the randomized benchmarking of transport, for different distances and trap depths, as specified above each subfigure. Top panel: atom survival for different move durations. The survival is fitted to a clipped Boltzmann distribution with linearly increasing temperature (Methods). The duration highlighted in blue is used for the data shown in the middle and bottom panels. Middle panel: coherence loss, comparing static and transported atoms. We plot the return probability to the initial quantum state after the randomized benchmarking sequence. The same single-qubit control sequence is applied in both cases, including the XY4 dynamical decoupling. Each point is the average result of 72 sequences of $N=80$ random single-qubit Clifford gates. The error bars are shown when larger than the marker size, and the curves are fits that include coherence and atom losses (Methods). Bottom panel: average instantaneous transport fidelity after a given number of moves. The fidelity is fitted from the return probability shown in the middle panel (Methods). The curve width represents the 68% confidence interval.
However, coherence-preserving transport between storage and adjacent interaction or readout zones might require covering large distances of $\sim$500 $\upmu$m. While moving atoms using acousto-optic deflectors (AODs) is now a well-established practice to resort them into a deterministic configuration barredo_atom-by-atom_2016, endres_atom-by-atom_2016 or to transport them coherently bluvstein_quantum_2022, shaw_multi-ensemble_2024, bluvstein_logical_2024, finkelstein_universal_2024, this distance is several times further than previously demonstrated single-atom transport bluvstein_quantum_2022, bluvstein_logical_2024. To address this challenge, we investigate the feasibility of coherent transport at this scale by loading atoms directly into tweezers made by AODs and performing the movement, given that pick-up and drop-off effects are not dependent on the distance of transport. We utilize the most wide-band commercially available AODs at near-infrared wavelengths, which for the optical parameters used here cover up to $500-600\ \upmu$m along one axis (Methods).
In probing the survival of atoms as they are moved long-distance with differing speeds, we find that the speed of transport is strongly constrained by cylindrical lensing, an effect that occurs as the AOD frequency is rapidly swept dickson_optical_1972, and which is more deleterious as the AOD field of view is scaled up (SI). In Fig. 5b, we present data that illustrate a striking consequence of this effect: with the same trap depth and trajectory, atoms can be transported considerably faster in a diagonal direction, utilizing two crossed AODs, when compared with using one AOD only. Such diagonal moves effectively turn cylindrical lensing into spherical lensing, thereby suppressing the worst impact. In the SI, we describe how diagonal motion can be used to transport atoms between zones without any penalty compared with a usual configuration relying on straight moves. We also discuss other known optical strategies that could completely remove any lensing friedman_acousto-optic_2000.
With diagonal moves, we first demonstrate in Fig. 5c the coherence of atoms being held in static tweezers in comparison with the coherence of atoms being transported by 610 $\upmu$m in 1.6 ms. We suppress any dephasing with one XY4 dynamical decoupling sequence per move, and do not observe loss in coherence between the static and transported case.
However, for any realistic application of coherent transport, atoms will have to be moved multiple times while maintaining coherence. Thus, we characterize the fidelity of the quantum channel defined by the coherent transport, through interleaved randomized benchmarking magesan_efficient_2012 (Fig. 5d). To our knowledge, this the first such rigorous transport channel fidelity characterization for neutral atoms. To maximize the dephasing cancellation, we apply dynamical decoupling in a transformed Clifford frame (Methods).
We perform this benchmarking technique for an average move distance from the storage zone to the interaction zone of $270\ \upmu$m (Fig. 5e,g), and also for a long move distance of $610\ \upmu$m (Fig. 5f). For each set of parameters (Fig. 5e-g), we first characterize the survival probability of an atom in a tweezer at the end of the sequence for different move durations (top panel). For selected durations, we then characterize the actual randomized benchmarking signal, the return probability to the initial quantum state (middle panel).
Notably, the resulting return probability data do not follow an exponential trend, because at large numbers of moves they become dominated by losses from the trap; as such, the fidelity for the transport channel depends on the number of previously executed moves. We thus define an instantaneous fidelity, the fidelity of the transport channel after a certain number of previous moves, shown in the bottom panel (Methods).
The instantaneous fidelity approaches a constant value for small number of moves ($\lesssim 30$), where losses are the sub-dominant error. This regime is most relevant for QEC, since one can in principle swap data qubits and ancilla qubits every few layers of gates chow_circuit-based_2024. We find that the fidelities in this regime approach $\sim$99.95for both transport distances of 270 $\upmu$m (Fig. 5e) and 610 $\upmu$m (Fig. 5g) in $1.0$ ms and in $1.6$ ms respectively using $0.28$ mK deep traps. We note that there is a small discrepancy between the early-time instantaneous fidelities obtained for 270 $\upmu$m and 610 $\upmu$m, and anticipate it could be remedied through more careful trap depth calibration along the trajectory.
For some applications, one might desire to optimize on the speed of movement and use a deeper trap to do so. In Fig. 5g, we show that atoms can be moved by 270 $\upmu$m in 400 $\upmu$s with a trap depth of $0.92$ mK, at the cost of a reduced fidelity of $\sim$99.8%. Comparing Fig. 5e and Fig. 5g illustrates an important trade-off pertaining to coherent transport: while atoms can be moved faster by increasing the trap depth $U$, the associated transport fidelity for small number of moves is also reduced. In the limit where noise is entirely induced by tweezer intensity fluctuations, this can be understood by noticing that the dephasing strength scales as $U$ when the required duration for long-distance transport merely scales as $U^{-1/2}$. We note that experimentally in static traps, we find an even stronger scaling of coherence time than linear in $U$, likely due to other sources of noise (Ext. Data Fig. 7g).
Finally, we envision utilizing multiple pairs of crossed AODs to cover the full extent of the array in Fig. 5a, with the demonstrated long-distance transport allowing overlap between adjacent AOD-pair controlled regions. With the layout presented here, four such regions would be necessary. Alternatively, additional scanning techniques, e.g., fast-scanning mirrors, can be used to position the field of view of single pair of crossed AODs across the full array in an iterative fashion.
As we lay out in detail in the SI, such techniques can be used also for initial rearrangement of atoms into the storage zone. For example, by implementing the recently developed “Tetris algorithm” wang_accelerating_2023 in four quadrants, with conservative estimates for relavant timings based on simulation, data, and previous experiments (SI Table I), we expect that we can sort the array in parallel in $\sim$137 ms or sequentially quadrant-by-quadrant in $\sim$522 ms. Further, this estimate is based on straight moves along a single AOD axis, which could be sped up by utilizing techniques for mitigating cylindrical lensing (SI §III.B).
Conclusion and outlook
We have shown that while scaling qubit numbers in optical tweezers to more than 6,100, we can simultaneously achieve record-high imaging survival alongside competitive imaging fidelity. We measure record room-temperature vacuum-limited lifetime and coherence time in alkali-metal atom tweezer arrays, with a high global single-qubit gate fidelity that is limited by technical noise. We additionally characterize the fidelity of a long-distance quantum transport channel utilizing randomized benchmarking and providing clear evidence for the potential to use this large scale array with a zone-based architecture for quantum computing (SI).
@@
Our results usher in a new generation of neutral atom quantum processors based on several thousand qubits, as described in the previous section. Additionally, large-scale programmable devices enabling advances in quantum metrology rosenband_exponential_2013, ludlow_optical_2015, norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024 and simulation haferkamp_closing_2020, orourke_entanglement_2023, julia-farre_amorphous_2024 are made accessible through this work. For example, our platform — with the demonstrated qubit numbers — could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2023. Tweezer clocks could be scaled using near-infrared, high-power tweezers for loading and imaging tao_high-fidelity_2024 before transferring atoms to magic-wavelength traps for clock operation norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024. We also foresee applications in quantum simulation for problems where boundary effects play an important role bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, orourke_entanglement_2023, which can be minimized with the large system sizes demonstrated here.
@@
Finally, our work indicates that further scaling of the optical tweezer array platform to tens of thousands of trapped atoms should be achievable with current technology, while essentially preserving high-fidelity control. In our present apparatus, several factors limit the number of sites. One limitation is the finite number of pixels of each SLM (which reduces the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers. By using available higher-resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently.
Furthermore, we observe worsening optical aberrations at tweezer powers greater than that in the present study due to thermal heating of the objective. This is the main limitation on atom number for the results in this work, even after aberrations were mitigated using the SLM (Methods). This constraint could be circumvented by utilizing an objective with a housing material that retains less heat or with integrated cooling strategies. Such upgrades should allow us to almost double the number of tweezers that we create using two fiber amplifiers. We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques to those utilized in this work (SI). Atom numbers may further be increased in our array with the same number of tweezers by utilizing enhanced loading brown_gray-molasses_2019 or re-loading techniques shaw_dark-state_2023, norcia_iterative_2024_mod. Already in the near term, we expect to increase the number of atomic qubits to over ten thousand with the current system using a subset of these techniques.